OpenStudy (anonymous):

find he zero for each function y=x^4-2x^3-x^2-2x

OpenStudy (anonymous):

y=x^4-2x^3-x^2-2x factor by grouping (x^4-2x^3)-(x^2-2x) x^3(x - 2) - x( x - 2) (x^3-x)(x-2) x(x^2-1)(x-2) x(x+1)(x-1)(x-2) x=0,1,-1,2 what i thought lol

OpenStudy (radar):

Did the sign of the -2x get changed in the solution. Looking at the third line at "-(x^2 -2x)" changes the sign of 2x It is a negative in the original equation.

OpenStudy (anonymous):

i didn't factor out "-x", if that's what you are going for. i chose to factor out a positive x to keep both factors of (x-2) on the left and (x-2) on the right equal.

OpenStudy (radar):

When the factors x(x+1)(x-1)(x-2) are multiplied the results give you x^4-2x^3-x^2+2x Check it out.

OpenStudy (anonymous):

even if you do pull out a negative x, and pull out a -x^2 likewise from the other side, the answer stays the same: y=x^4-2x^3-x^2-2x -x^3(2-x) + x (2 - x) (x-x^3)(2-x) -x(x^2-1)(2-x) x(x^2-1)(x+2) x(x+1)(x-1)(x+2) x=0,-1,1,-2

OpenStudy (anonymous):

hmm lemme see

OpenStudy (anonymous):

i got a different answer that second time actually - (x+2), which is the needed term to make the -2x. (x-2)(x^2 - 1) = neg*neg = pos 2x (x+2)(x^2 - 1) = pos*neg = neg 2x apparently i dont check my work close enough lol. thanks for looking at that lol !

OpenStudy (radar):

Is 1 a solution? Does it result in a 0? The -2 and 0 are solutions.

OpenStudy (radar):

I'm checking the -1 now for a zero.

OpenStudy (radar):

The -1 solution results in 4.

OpenStudy (radar):

To be honest, I think the problem was meant to be a +2x for the last term. lol

OpenStudy (radar):

@cherry33, do you have a solution sheet?

OpenStudy (radar):

And have you copies the equation correctly?